15.01 CASH CONVERSION CYCLE

Parramore Corp has $20 million of sales, $2 million of inventories, $4 million of receivables, and $3 million of payables. Its cost of goods sold is 80% of sales, and it finances working capital with bank loans at an 8% rate. Assume 365 days in year for your calculations. Do not round intermediate steps.

  1. What is Parramore's cash conversion cycle (CCC)? Do not round intermediate calculations. Round your answer to two decimal places.
    _____ days
  2. If Parramore could lower its inventories and receivables by 10% each and increase its payables by 10%, all without affecting sales or cost of goods sold, what would be the new CCC? Do not round intermediate calculations. Round your answer to two decimal places.
    ______ days
  3. How much cash would be freed up, if Parramore could lower its inventories and receivables by 10% each and increase its payables by 10%, all without affecting sales or cost of goods sold? Do not round intermediate calculations. Round your answer to the nearest cent. Write out your answer completely. For Example, 13.2 million should be entered as 13,200,000.
    $ ______
  4. By how much would pretax profits change, if Parramore could lower its inventories and receivables by 10% each and increase its payables by 10%, all without affecting sales or cost of goods sold? Do not round intermediate calculations. Round your answer to the nearest cent. Write out your answer completely. For Example, 13.2 million should be entered as 13,200,000.
    $ ______
Answer #1

Sales = 20 Million

Inventory = 2 million

Receivables = 4 million

Payables = 3 million

COGS = 80%* Sale = 0.80*20 = 16 million

Cash Conversion Cycle = Days Sales outstanding + Days Inventory Outstanding - Days Payables Outstanding

= 73 + 45.63 - 68.44

= 50.19 Days

Days Sales Outstanding = 365/Receivable Turnover = (365*Account Receivable)/Sales

= 365 * 4 million/ 20 million

= 73 days

Days Inventory Outstanding = 365/Inventory Turnover = (365*Inventory)/Cost of Goods Sold

= 365* 2 million / 16 million

= 45.63 days

Days Payables Outstanding = 365/Payables Turnover = (365*Account Payable)/Cost of Goods Sold

= 365* 3 Million / 16 Million

= 68.44 days

If inventories are lowered by 10%, then new inventory level = 2million*0.9 = 1.8 million

If receivables are lowered by 10%, then new receivable level = 4million *0.9 = 3.6 million

If payables increase by 10% = $3million x 1.1 = $3.3million

Days Sales Outstanding = 365/Receivable Turnover = (365*Account Receivable)/Sales

= 365 * 3.6 million/ 20 million

= 65.70 days

Days Inventory Outstanding = 365/Inventory Turnover = (365*Inventory)/Cost of Goods Sold

= 365* 1.8 million / 16 million

= 41.06 days

Days Payables Outstanding = 365/Payables Turnover = (365*Account Payable)/Cost of Goods Sold

= 365* 3.3 Million / 16 Million

= 75.28 days

Cash Conversion Cycle = Days Sales outstanding + Days Inventory Outstanding - Days Payables Outstanding

= 65.70 + 41.06 - 75.28

= 182.88 Days

Cash freed up:

Inventory = (45.63 – 41.06) * 43835.62 = $20,0328.8

Receivables = (73 – 65.7) * 54794.52 = $400000

Payables = (68.44 – 75.8) * 43835.62 = $324384

Cash freed up = $200328.8 + $400000 – $ 324384

= $275945

Increase in pre-tax profit = $275945 * 0.08 = $22075.6